OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 14x + 18y = -54 -4x – 14y = 42

OpenStudy (anonymous):

multiply the first equation across by 2 multiply the 2nd equation across by 7 28x + 36y = -108 -28x -98y = 294 ------------------ 0x - 62y = 186 y = -186/62 =-3 plug y = -3 into one of the equations: -4x - 14(-3) = 42 -4x +42 = 42 x = 0

hero (hero):

y = (-54-14x)/18 y = (42+4x)/-14 y = y (-54-14x)/18 = (42+4x)/-14 -14(-54-14x) = 18(42+4x) 756 + 196x = 756 + 72x 124x = 0 x = 0 y = -3

hero (hero):

My steps probably make no sense whatsoever.

hero (hero):

I set y = y to eliminate y as a variable.

hero (hero):

Then I cross multiplied and solved for x. Then used x = 0 to find y.

OpenStudy (anonymous):

The problem asked for elimination method, so ...

hero (hero):

My approach is valid for elimination method

OpenStudy (anonymous):

yeah true. my bad

OpenStudy (anonymous):

conheomap-- at the end of your problem how did you get 42-4x+42=42x=0 ?? wouldn't the 42s cancel out, and you would be left with -4x?? I have this same problem that I am trying to work out the answer to it right now, but it doesn't seem to be going right.. HELP!